• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 15:48:57 on February 28, 2011

    Entry number 349215

    keyword=Update of HRS detectors and Create a new data base ~/DB/20110228/ due to move cables of S1-6R, S2-8R

    It looks like all detectors in L-HRS are fixed. For R-HRS, only PS-18R is
    still noisy. I traced the signal from upstairs patch panels and
    downstairs pp, and they are clean. I changed its BNCtoECL black cable
    from PP to ADC and also moved it to other ADC channel, but the broad
    pedestal peak is still here. I really don't know which part has problem now.
    I moved the cables of S1-6R,S2-8R at patch panel upstairs to the last two
    inputs of S1-ADC group (after S0A,S0B), and I created a new data base
    ~/DB/20110228 to keep this change.

    A copy of this log entry has been emailed to: danez@jlab.org